1. A winch lifted a mass of 2500 kg through a height of 22 m in 15 s. Calculate the power in
hp.

2. A 50 kg is raised vertically 8 m. What is the potential energy?
1
Expert's answer
2020-08-06T16:45:09-0400

1. The power is (by definition):


"P = Fv"

where "F = mg" is the gravity force against which the winch lifted a mass, "v = h\/t" is the speed of this mass ("h =22m, t=15s" ). Thus, obtain:


"P = \dfrac{mgh}{t} = \dfrac{2500\cdot9.81\cdot 22}{15} = 35970\space W"

In order to convert it to hp, one should multiply it by 0.00134:


"P = 35970\cdot0.00134 \approx 48.2\space hp"

2. The potential energy can be caculated with the following formula:


"U = mgh = 50kg\cdot 9.81 m/s^2\cdot 8m = 3924J"

Answer. 1) 48.2 hp, 2) 3924 J.